A world-class sprinter can burst out of the blocks to essentially top speed (of about ) 11.0 m/s in the first 10.9 m of the race. What is the average acceleration of this sprinter in m/s2 ?
1
Expert's answer
2012-10-08T09:52:20-0400
Let's find the time needed a sprinter to reach his top speed. We know that

V = at ==> a = V/t

and

D = at²/2 = Vt/2 ==> t = 2D/V.

Therefore

a = V/(2D/V) = v²/(2D) = (11.0[m/s])²/(2*10.9[m]) ≈ 5.55 m/s².

So, the average acceleration of this sprinter is about 5.55 m/s².